OpenStudy (anonymous):

find the slope of the tangent line to f(x) at the point where x=0 f(x)= cosh(3x) + sinh(4x)

OpenStudy (anonymous):

f(x)= cosh(3x) + sinh(4x) f'(x)= 3sinh(3x) + 4cosh(4x) f'(x)= 3sinh(3x) + 4cosh(4x)]x=0 f'(0)= 3sinh(0) + 4cosh(0)=4

OpenStudy (anonymous):

delarousse any question?

OpenStudy (anonymous):

thank you